# Numerical investigation of buoyancy-aided mixed convective flow past a square cylinder inclined at 45 degrees

### v1

@@ III.3 Time-averaged flow
Richardson number (Ri)
Drag coefficient
($C_{D}$)
RMS lift coefficient ($C_{L,rms}$) Mean Nusselt number ($\overline{Nu}$)
Mean ($\overline{C}_{D}$) RMS ($C_{D,rms}$)
@@
The variation of the integral parameters with Ri is shown in Table 8. $\overline{C}_{D}$ and $\overline{Nu}$ show an increase with increasing Ri. The buoyancy-driven flow in the vicinity of the cylinder increases both pressure drag and skin friction drag on the cylinder. On the other hand, as Ri increases, RMS quantities decrease. All quantities presented in Table 8 are near-field quantities, and hence, when the near-field is stabilized, these values become constant with no periodicity. Hence, RMS quantities fall to zero beyond the critical Ri. The RMS Nusselt number has not been reported as it is extremely small for all Ri. The mean Nusselt number presents an interesting variation. Below Ri $=0.7$, the Nusselt number increases monontonically, and it decreases to a local minimum value at Ri $=0.7$, and increases subsequently for Ri $>$ 0.7. Below 0.7, flow acceleration due to buoyancy forces causes thinning of the thermal boundary layer, leading to higher Nusselt numbers for the entire cylinder. There is a greater increase in the Nusselt number of the flow-facing surfaces compared to the decrease in the Nusselt number of the wake-facing surfaces, leading to an overall increase in the Nusselt number of the cylinder. However, at Ri $=0.7$, as discussed earlier, the Nu of the wake-facing surfaces sharply decreases due to the inhibition of vortex shedding, and this decrease is greater than the increase in the Nusselt number of flow-facing surfaces, leading to an overall decrease in the Nusselt number of the cylinder. Beyond Ri $=0.7$, an increase in cylinder Nu is seen, similar to the trend below Ri $=0.7$.
@@
Figure 15 reveals the variation of recirculation and inversion lengths with Ri. For calculation of the inversion length, time-averaged vorticity data is taken for all streamwise locations at a transverse location X $=0.1$. The location of the first sign change in the vorticity data yields the inversion length. To ensure independence of the result from the choice of transverse location, we have also calculated inversion length using X $=0.2$, and the results reveal no variation based on the choice of X. For the calculation of recirculation length, time-averaged streamwise velocity data along the centerline is considered, and the location of the sign change in this data yields the recirculation length. There is a monotonic decrease in the inversion length up to Ri $=0.6$, followed by a slight increase at Ri $=0.7$ and a subsequent decrease for higher Ri. The decrease in inversion length is due to the increasing strength of the natural convective vorticity and the diminishing strength of the forced convective vorticity with increasing Ri. At low Ri, the forced convective component is strong, but it diminishes with downstream distance, while the natural convective component is weak in the near-field, but its strength grows with downstream distance. When Ri is increased, the strength of the natural convective vorticity closer to the cylinder also grows, as the buoyancy-driven flow transforms the entire flow field into a plume, while the wake behavior weakens. The recirculation length shows an interesting variation, which was discussed briefly with Figure 13(a). The recirculation length increases up to Ri $=0.7$ where it peaks, followed by a decrease. Prior to Ri $=0.7$, vortex shedding actively occurs in the flow. Due to the buoyancy, there is an inhibited interaction of the two shear layers as seen in near-field plots of vorticity and temperature (see Figures 4 and 6). This increases the shear layer length and consequently, the recirculation length, which increases up to Ri $=0.7$, where there is no vortex shedding and near-field flow becomes steady. Since the flow is steady, there is no mutual interaction between the opposite shear layers, which form a steady separation bubble, and the effect of increasing buoyancy is to cancel the vorticity caused by the forced convective part, which reduces the length of the separation bubble.
@@
Streamwise velocity, transverse velocity, vorticity, and temperature profiles at different streamwise locations for different Ri are shown in Figures 16, 17, 18 and 19, respectively. Figure 16 shows the deficit of momentum in the near-field associated with wake-like behavior and a recovery and subsequent excess of momentum in the intermediate and far-fields. In the near-field, the deficit of momentum is seen for all Ri, resembling the profile seen in wakes. With increased downstream distance, the recovery of momentum occurs, starting earlier downstream for higher Ri, followed by lower Ri later downstream. At Y $=30$, the recovery of momentum is complete for all positive Ri, and there is an excess of momentum seen for higher Ri, which grows more pronounced further downstream, resembling the profile seen in jets. Therefore, the flow field downstream starts out as that of a wake at low Ri and slowly transforms into that of a plume as Ri is increased, first far downstream, then near the cylinder as well. Such behavior is attributed to the buoyancy-driven flow, whose tendency is to cause a jetting effect in the complete flow field, starting from the far downstream region and slowly coming towards the cylinder. Figure 17 shows the variation of transverse velocity, which is positive in the left half and negative in the right half of the flow for all positive Ri at all downstream locations. Figure 18 reveals the inversion of vorticity in the near and intermediate fields. For higher Ri, the inversion occurs earlier downstream, while for lower Ri, it occurs further downstream as the strength of the natural convective vorticity is lesser at lower Ri. The peak vorticity is seen to be higher for higher Ri. Figure 19 shows the temperature profiles for different streamwise locations. Overall, the temperature shows a dissipating trend with increased streamwise distance. We also observe that prior to the onset distance of the far-field unsteadiness, the temperature profile width for the cases above the critical Ri is much lower than that of the cases below the critical Ri (see Figure 19(e)). However, after the onset distance of the far-field unsteadiness, the profile widens drastically, and it is comparable with the cases below the critical Ri (see Figure 19(f)).
The inversion of vorticity has been reported in the literature for isothermal confined flows. [13, 14, 15, 16] The interplay between several sources of vorticity in the flow causes the reversal of vorticity in the wake of the cylinder. In unconfined flows, a reversal of the Bénard-von Kármán vortex street takes place when the obstacle (bluff body or aerofoil) is imparted a periodic motion. Under appropriate conditions of the shape of the body or the control parameter, the shed vortices cross the centerline and switch their position after traveling downstream, which manifests as a jet profile with a momentum excess in the time-averaged velocity plot, compared to the original wake profile with a momentum deficit. [40, 41, 42] We have seen similar behavior in the flow structures obtained for our case (see Figure 3), however, the shed vortices do not physically convect to the opposite side in our case. Rather, there is an inversion of vorticity due to the interplay between the forced convective and natural convective vorticity, as discussed earlier.
The forced convective component of the flow causes shear layer separation at the downstream corners of the cylinder, causing a deficit of momentum in the wake. On the other hand, the natural convective component causes jet-like behavior downstream, competing with the deficit caused by the wake, with the respective vorticities signed oppositely to each other. The opposing effects of these two flow components lead to a vorticity inversion. This inversion does not occur for Ri $=0.0$ due to the absence of the buoyancy-driven flow at this Ri. For positive Ri, vorticity inversion occurs downstream at an inception point, which comes further upstream as Ri is increased (see Figure 15), due to the increased buoyancy (see Figures 3 and 9). Prior to the reversal of vorticity, the downstream region has a wake-like behavior with a momentum deficit, and beyond the reversal of vorticity, the dowstream region exhibits a plume-like behavior, with momentum excess seen in the time-averaged velocity profile (see Figure 16). The strength of the wake decreases with increased downstream distance, while the plume shows the opposite behavior. Such behavior resembles that of fish-like swimming and flapping foil propulsion in the time-averaged sense, but not in the instantaneous sense, because the shed vortices do not physically travel to the opposite side in our case. Instead, the greater strength of the natural convective vorticity causes the change of sign of vorticity on either side of the centerline by dominating over the forced convective vorticity, whose strength diminishes with increased downstream distance. The far-field unsteadiness also shows characteristics dissimilar to flapping foil cases, in that there are no discrete vortices in the far-field of the present case. The shear layers that were already shed in the near-field suddenly undergo large-scale undulation in the far-field, which appears to be due to the pronounced strength of the natural convective jet.
@@
III.4 Momentum deficit and vortex shedding suppression in the near-field
@@
As discussed earlier, shear layer separation occurs at the downstream corners of the cylinder, leading to a deficit of velocity in the cylinder wake. Beyond a critical Re, the flow transitions to unsteadiness, and the Bénard-von Kármán vortex street is seen. However, when the cylinder is heated, density differences cause buoyancy-driven momentum addition, which causes flow acceleration in the wake, particularly around the centerline. This leads to the suppression of vortex shedding in a similar mechanism as other flow control techniques, which prevent the interaction of the two shear layers that originate at the downstream corners.
@@
The increased momentum in the wake at higher Ri suggests that the effect of buoyancy is to add a source that effectively behaves as a jet without the addition of any new momentum, as explained below. At lower Ri, the effective jet caused by buoyancy does not possess sufficient momentum to affect the interaction between the two shear layers. As Ri is increased, a larger velocity of the effective jet causes the undulation of the oppositely signed shear layers to decrease. The interaction between the shear layers is affected by the effective jet, which behaves in a similar way as a splitter plate [43] or a blowing jet. [44] Below the critical Ri, this manifests as an increase in the recirculation length (see Figure 15) due to the increased shear layer length prior to roll-up and detachment. Additionally, the cancellation of the forced convective vorticity by the opposite natural convective vorticity inhibits the interaction between the shear layers. The hindered interaction between the shear layers leads to a loss of strength due to the inability to mutually gain further circulation, leading to a situation where detachment of the shear layers is not possible [2] and the vortices remain as a steady separation bubble with no roll-up, beyond the critical Ri of 0.68 (see Figure 4).
”Momentum deficit” or ”momentum addition” may be misunderstood as the disappearance or appearance of momentum in the domain. It is important to note that there is no disappearance or appearance of momentum in the domain. The shear layer separation at the lateral corners of the cylinder causes the flow to accelerate around the cylinder, while the flow that strikes the cylinder head-on is decelerated, leading to a large momentum deficit in the wake. Beyond the cylinder, the flow velocity near the centerline is low, while the flow velocity parallel to the centerline is high. When the cylinder is heated, the buoyancy-driven flow behaves like a jet in the downstream region, accelerating the flow around the centerline and decelerating the flow parallel to the centerline. These flow accelerations are nominally referred to as ”momentum deficit” and ”momentum addition,” respectively.
III.5 Unsteadiness of the far-field
@@
In previous studies, the far-field unsteadiness has not been captured due to the usage of a streamwise domain length less than the onset distance of the far-field unsteadiness, resulting in an inaccurate prediction of steady flow in the entire domain. To avoid such issues and to fully capture the physics of the flow field, we have considered a downstream length ($L_{d}$) of 110d. Instantaneous and time-averaged contours in Figures 5, 3, 11 and 9 reveal that at Ri $=0.0$, the far-field is steady, while for Ri $>0.0$, the far-field is unsteady, which can be attributed to the increasing strength of the natural convective flow which makes the far-field behave like a plume. At the critical Ri and beyond, the far-field develops large-scale unsteadiness that is very similar to plume-like unsteadiness reported in the literature. [24, 25, 26, 31] This occurs in conjunction with the suppression of near-field vortex shedding. There is no range of Ri for which the flow field is entirely steady. As Ri is increased beyond the critical Ri, the point of the onset of the far-field unsteadiness appears to come closer to the cylinder, which further indicates that the plume strength increases closer to the cylinder on increasing Ri. For Ri $\leq 0.6$, there is a widening of the plume in the intermediate-field itself, and in the far-field, the plume narrows (see Figures 11 and 9). For Ri $\geq 0.7$, the plume is narrow in the near and intermediate fields, and it sharply widens in the far-field with the onset of the far-field unsteadiness. The natural convective component grows stronger with increased downstream distance, and the opposite occurs with the forced convective component. Due to this, there is a large momentum excess far downstream (Y $>$ 60) which causes strong jet-like behavior in the far-field. This can be seen in Figures 16 and 13(a). The jetting effect in the far-field is responsible for the far-field plume-like unsteadiness, which shows similar characteristics as the instabilities observed in natural convection plumes. [24, 25, 26, 31] The flow structures observed in the far-field require further detailed investigation for a complete understanding of the flow physics in the far-field.
The observed transition sequence is in contrast with that reported by Dushe [17] for the $\alpha=0^{\circ}$ case. For $\alpha=0^{\circ}$, there is a range (0.15-0.58) where the entire flow field is steady in the near-field as well as the far-field. However, for our $\alpha=45^{\circ}$ case, the flow immediately transitions from near-field unsteadiness to far-field unsteadiness, with no intermediate Ri at which there is steady flow in the entire domain. At any positive Ri, the flow is unsteady in at least one region of the domain.

### v3

@@ III.2 Time-Averaged Flow
Richardson number (Ri) Drag coefficient ($C_{D}$) RMS lift coefficient ($C_{L,\text{rms}}$) Mean Nusselt number ($\overline{Nu}$)
Mean ($\overline{C}_{D}$) RMS ($C_{D,\text{rms}}$)
@@
The variation of the integral parameters with Ri is listed in Table 7. $\overline{C}_{D}$ and $\overline{Nu}$ show an increase with increasing Ri. The buoyancy in the vicinity of the cylinder increases both pressure drag, and skin friction drag on the cylinder. On the other hand, as Ri increases, RMS quantities decrease as expected. All quantities presented in Table 7 are near-field quantities, and hence, when the near-field is stabilized, these values become constant with no periodicity. Hence, RMS quantities fall to zero beyond the critical Ri. The RMS Nusselt number has not been reported as it is extremely small for all Ri. The mean Nusselt number presents an interesting variation. Below Ri $=0.7$, the Nusselt number increases monotonically, decreasing to a local minimum value at Ri $=0.7$ and subsequently increasing for Ri $>$ 0.7. Below 0.7, flow acceleration due to buoyancy forces causes thinning of the thermal boundary layer, leading to higher Nusselt numbers for the entire cylinder. There is a greater increase in the Nusselt number of the flow-facing surfaces compared to the decrease in the Nusselt number of the wake-facing surfaces, leading to an overall increase in the Nusselt number of the cylinder. However, at Ri $=0.7$, as discussed earlier, the Nu of the wake-facing surfaces sharply decreases due to the inhibition of vortex shedding, and this decrease is greater than the increase in the Nusselt number of flow-facing surfaces, leading to an overall decrease in the Nusselt number of the cylinder. Beyond Ri $=0.7$, an increase in cylinder Nu is seen, similar to the trend below Ri $=0.7$.
@@
Figure 15 reveals the variation of recirculation and inversion lengths with Ri. For calculation of the inversion length, time-averaged vorticity data is taken for all streamwise locations at a transverse location X $=0.1$. The location of the first sign change in the vorticity data yields the inversion length. To ensure independence of the result from the choice of transverse location, we have also calculated inversion length using X $=0.2$, and the results reveal no variation based on the choice of X. For the calculation of recirculation length, time-averaged streamwise velocity data along the centerline is considered. The location of the sign change in this data yields the recirculation length. There is a monotonic decrease in the inversion length up to Ri $=0.6$, followed by a marginal increase at Ri $=0.7$ and a subsequent reduction for higher Ri. The decrease in inversion length is due to the increasing strength of the natural convective vorticity over the strength of the forced convective vorticity with increasing Ri. The forced convective component is strong at low Ri but diminishes with downstream distance. At the same time, the natural convective component is weak in the near-field, but its strength grows with downstream distance. When Ri is increased, the strength of the natural convective vorticity closer to the cylinder also grows. The recirculation length shows an interesting variation, which was discussed briefly with Figure 13(a). It increases up to Ri $=0.7$, where it peaks, followed by a decrease. Prior to Ri $=0.7$, vortex shedding actively occurs in the flow. Due to the jet formed by buoyancy, the interaction between the two shear layers is inhibited, as seen in near-field plots of vorticity and temperature (see Figures 4 and 7). This increases the shear layer length and, consequently, the recirculation length, which increases up to Ri $=0.7$. Beyond Ri $=0.7$, there is no vortex shedding, and the near-field flow becomes steady. Since the flow is steady, there is no mutual interaction between the opposite shear layers, which form a steady separation bubble, and the effect of increasing buoyancy beyond the critical suppression value is to add momentum to the separation bubble, reducing its length.
@@
Streamwise velocity, transverse velocity, vorticity, and temperature profiles at different streamwise locations for different Ri are shown in Figures 16, 17, 18 and 19, respectively. Figure 16 shows the deficit of momentum in the near-field associated with wake-like behavior and a recovery and subsequent excess of momentum in the intermediate and far-fields. In the near-field, the momentum deficit is seen for all Ri, resembling the profile seen in wakes. With increased downstream distance, the recovery of momentum occurs, starting earlier downstream for higher Ri, followed by lower Ri later downstream. At Y $=30$, the recovery of momentum is complete for all positive Ri, and there is an excess of momentum seen for higher Ri, which grows more pronounced further downstream, resembling the profile seen in jets. Therefore, the flow field downstream starts resembling a wake at low Ri and slowly transforms into that of a plume as Ri is increased, first far downstream, then near the cylinder. Such behavior is attributed to buoyancy, which tends to cause a jetting effect in the entire flow field, starting from the far downstream region and slowly moving towards the cylinder. Figure 17 shows the variation of transverse velocity, which is positive in the left half and negative in the right half of the flow for all positive Ri at all downstream locations. Figure 18 reveals the inversion of vorticity in the near and intermediate fields. For higher Ri, the inversion occurs earlier downstream, while for lower Ri, it occurs further downstream as the strength of the natural convective vorticity is lesser at lower Ri. The peak vorticity is seen to be higher for higher Ri. Figure 19 shows the temperature profiles for different streamwise locations. Overall, the temperature shows a dissipating trend with increased streamwise distance. We have also observed that prior to the onset distance of the far-field unsteadiness, the temperature profile width for the cases above the critical Ri is much lower than that of the cases below the critical Ri (see Figure 19(e)). However, after the onset distance of the far-field unsteadiness, the profile widens drastically, and it is comparable with the cases below the critical Ri (see Figure 19(f)).
@@
III.3 Momentum Deficit and Vortex Shedding Suppression in the Near-Field
@@
As discussed earlier, shear layer separation occurs at the lateral corners/edges of the cylinder, leading to a deficit of velocity in the cylinder wake. Beyond a critical Re, the flow transitions to unsteadiness, and the Bénard-von Kármán vortex street is seen. However, when the cylinder is heated, density differences cause momentum addition, leading to flow acceleration in the wake, particularly around the centerline. This leads to the suppression of vortex shedding in a mechanism similar to other flow control techniques, which prevents the interaction of the two shear layers originating at the two lateral corners.
@@
The increased momentum in the wake at higher Ri suggests that the effect of buoyancy is to add a source that effectively behaves similarly to a jet, thus adding equivalent momentum to the wake, as explained below. At lower Ri, the effective jet caused by buoyancy does not possess sufficient momentum to affect the interaction between the two shear layers. As Ri is increased, a larger velocity of the effective jet causes the undulation of the oppositely signed shear layers to decrease. The interaction between the shear layers is affected by the effective jet, which behaves in a similar way as a splitter plate [37] or a blowing jet [38]. Below the critical Ri, this manifests as an increase in the recirculation length (see Figure 15) due to the increased shear layer length prior to roll-up and detachment. Additionally, the cancellation of the forced convective vorticity by the opposite natural convective vorticity inhibits the interaction between the shear layers. The hindered interaction between the shear layers leads to a loss of strength due to the inability to gain further circulation mutually, leading to a situation where detachment of the shear layers is not possible [2] and the vortices remain as a steady separation bubble with no roll-up beyond the critical Ri (see Figure 4).
III.4 Unsteadiness of the Far-Field
The effect of the chosen domain in mixed convective flows is significant and changes the observed physics completely. In previous works by Sharma and Eswaran [21], Hasan and Ali [39], and others, the streamwise length of the domain has not exceeded 50 times the projected length of the cylinder. Due to this, transitions to unsteadiness in the far field much beyond the cylinder have not been captured, and the flow field has been remarked as steady, although it may not be the case far downstream of the cylinder. To fully capture the physics of the flow field, we have considered a downstream length ($L_{d}$) of $110d$. It must be noted that the results contained in the present work pertain only to the domain size chosen here. Additional physics might be seen with the choice of an even larger streamwise domain.
@@ III.4 Unsteadiness of the Far-Field
Instantaneous and time-averaged contours in Figures 6, 3, 11 and 9 reveal that at Ri $=0.0$, the far-field is steady, while for Ri $>0.0$, the far-field is unsteady, which can be attributed to the increasing strength of the natural convective flow which makes the far-field behave like a plume. At the critical Ri and beyond, the far-field develops large-scale unsteadiness that is very similar to plume-like unsteadiness reported in the literature [25, 26, 27, 31]. This occurs in conjunction with the suppression of near-field vortex shedding. There is no range of Ri for which the flow field is entirely steady. As Ri is increased beyond the critical Ri, the point of onset of the far-field instability appears to come closer to the cylinder, which further indicates that the plume strength increases closer to the cylinder with increasing Ri. For Ri $\leq 0.6$, there is a widening of the plume in the intermediate-field itself, and in the far-field, the plume narrows (see Figures 11 and 9). For Ri $\geq 0.7$, the plume is narrow in the near and intermediate fields, and it sharply widens in the far-field with the onset of the far-field unsteadiness. The natural convective component grows stronger with increased downstream distance, and the opposite phenomenon occurs with the forced convective component. Due to this, there is a large momentum excess far downstream (Y $>$ 60), which causes strong jet-like behavior in the far-field. This can be seen in Figures 16 and 13(a). The jetting effect in the far-field is responsible for the plume-like unsteadiness of the far-field, which shows similar characteristics as the instabilities observed in natural convection plumes [25, 26, 27, 31]. The flow structures observed in the far-field require further detailed investigation to gain a complete understanding of flow physics.
The critical Ri for the transition to steady flow for both square and circular cylinders is 0.15, as reported by Sharma and Eswaran [21], and Chang and Sa [22]. However, the critical Ri is between 0.6 and 0.7 for our inclined square cylinder, which is significantly greater than that reported for square and circular cylinder cases. This large variation may be attributed to the (i) thickness and vorticity of the separation shear layer contained in it and/or (ii) afterbody effect, that is, the part of the obstacle beyond the separation of the flow from it. The shape of the obstacle and the corresponding afterbody are expected to affect the transition sequence in mixed convective flows past obstacles. The exact details of the effect require further investigation.
Figure 20: Variation of (a) peak total vorticity, and (b) peak baroclinic vorticity in the streamwise direction, for various Ri. The peak vorticity is the magnitude of the maximum vorticity at any streamwise location.
The curl of the momentum equation (refer Equation 2) results in the vorticity transport equation and is obtained as:
[…]
The second term on the right side of Equation 10 is the baroclinic vorticity production term due to the buoyancy effect. Buoyancy causes density gradients to arise in the fluid. The baroclinic vorticity appears as a source term in the vorticity transport equation when the density and pressure gradients are not aligned in a fluid. It is seen that this vorticity plays a key role in the suppression of vortex shedding as well as the far-field unsteadiness.
Figure 4 shows the total vorticity (flood) and the baroclinic vorticity (lines) in the near-field region. At Ri $=0.0$, there is no baroclinic vorticity due to the absence of thermal gradients. However, for positive Ri, baroclinic vorticity affects the near-field flow structures. The shed vortices appear to be stretched in the transverse direction, matching the shape of the baroclinic vortices. This effect is intensified at higher Ri. On comparing the shed vortices at Ri $=0.7$ with those at Ri $=0.0$, it is clear that the vortices at higher Ri are thinner and elongated compared to those at lower Ri. For Ri $<0.7$, the baroclinic vorticity follows an alternating pattern, with vortices of opposite signs being shed one after the other. The regions with a high magnitude of total vorticity are accompanied by positive baroclinic vorticity, while regions with a low magnitude of total vorticity are accompanied by negative baroclinic vorticity. This reveals that baroclinic vorticity increases the vorticity content in high-vorticity regions while cancelling the vorticity in the low-vorticity regions.
Figure 5 shows the total vorticity (flood) and the baroclinic vorticity (lines) in the far-field region. It is seen that the baroclinic vorticity term plays a significant role in the development of the far-field unsteadiness. The baroclinic vorticity at Ri $=0.2$ creates unsteady flow structures in the far-field, unlike the forced convective flow case. It can also be seen that at this Ri, the effect of the baroclinic vorticity diminishes with increased downstream distance, as does the total vorticity. At Ri $=0.4$, the baroclinic vorticity is more pronounced. Regions of positive baroclinic vorticity increase the magnitude of total vorticity (in both positive and negative vorticity streams), while regions of negative baroclinic vorticity decrease the magnitude of total vorticity in both vorticity streams. A similar trend is observed for Ri $=0.6$. However, the strength of baroclinic vorticity increases for Ri beyond 0.7, which causes large-scale undulation of the shear layer, leading to unsteady flow structures. This effect is intensified at higher Ri. Similar to Ri $\leq$ 0.7, the regions of positive baroclinic vorticity enhance the magnitude of total vorticity, while regions of negative baroclinic vorticity minimize or cancel the magnitude of total vorticity. This behaviour is similar to that seen in the near-field.
Figure 20 shows the streamwise variation of peak total vorticity and peak baroclinic vorticity for different Ri. For Ri $<0.7$, the total vorticity shows a decaying trend in the streamwise direction due to the diffusion of vorticity as it is convected away from the cylinder. Above Ri $=0.7$, the vorticity decays sharply after the cylinder and increases immediately thereafter, remaining constant until the onset of the far-field unsteadiness around Y $=65$. Both Ri of 0.7 and 1.0 show a maxima in the total vorticity with the onset of the far-field unsteadiness. This reveals that the shedding of vortices in the far-field occurs when the shear layer is unable to contain further vorticity beyond a prescribed limit. After the shedding begins, the vorticity once again decays with space due to the diffusion effect. On examining the plot of peak baroclinic vorticity, it is evident that a behavior similar to total vorticity is shown, with maxima in baroclinic vorticity visible in the far-field region.